# second generation 3G third generation

### v1

@@
Gilderlan Tavares de Araújo, , André L. F. de Almeida, , Bruno Sokal, , Gabor Fodor, , Paulo R. B Gomes The authors thank the partial support of FUNCAP under grant ITR-0214-00041.01.00/23, and the National Institute of Science and Technology (INCT-Signals) sponsored by Brazil’s National Council for Scientific and Technological Development (CNPq) under grant 406517/2022-3.Gilderlan T. de Araújo and Paulo. R. B. Gomes are with the Federal Institute of Ceará, e-mails: {gilerlan.tavares,gomes.paulo}@ifce.edu.br.André L. F. de Almeida and Bruno Sokal are with the Teleinformatics Department, Federal University of Ceará, Fortaleza-CE, e-mail: {andre,brunosokal}@gtel.ufc.br.G. Fodor was funded by the Swedish Strategic Foundation Future Software Systems Project SAICOM, grant no: FUS21-0004
@@ I Introduction
$\textrm{vec}(\boldsymbol{ABC})=(\boldsymbol{C}^{\textrm{T}}\otimes\boldsymbol{A})\textrm{vec}(\boldsymbol{B}).$ (1)
$\textrm{diag}(\boldsymbol{a})\boldsymbol{b}=\textrm{diag}(\boldsymbol{b})\boldsymbol{a}.$ (2)
If $\boldsymbol{B}$ is a diagonal matrix, we have:
$\textrm{vec}(\boldsymbol{ABC})=(\boldsymbol{C}^{\textrm{T}}\diamond\boldsymbol{A})\textrm{vecd}(\boldsymbol{B}).$ (3)
@@ II System Model
$\boldsymbol{y}_{t}=\sum_{q=1}^{Q}\boldsymbol{H}^{(q)}\boldsymbol{S}^{(q)}\boldsymbol{G}^{(q)}\boldsymbol{x}_{t}+\boldsymbol{z}_{t}^{(q)},$ (4)
where $\boldsymbol{H}^{(q)}\in\mathbb{C}^{M_{R}\times\overline{N}}$ and $\boldsymbol{G}^{(q)}\in\mathbb{C}^{\overline{N}\times M_{T}}$ represent the RIS-BS and UT-RIS channels, respectively, while $\boldsymbol{S}^{(q)}\in\mathbb{C}^{\overline{N}\times\overline{N}}$ denotes the (non-diagonal) scattering matrix associated with the $q$-th group. Additionally, $\boldsymbol{x}_{t}\in\mathbb{C}^{M_{T}\times 1}$ represents the data symbol vector, while $\boldsymbol{z}_{t}^{(q)}$ corresponds to the additive white Gaussian noise (AWGN) term.
@@
$\boldsymbol{y}_{i,k,t}=\sum_{q=1}^{Q}\boldsymbol{H}^{(q)}\boldsymbol{S}_{k}^{(q)}\boldsymbol{G}^{(q)}_{i}\text{diag}(\boldsymbol{w}_{k})\boldsymbol{x}_{t}+\boldsymbol{z}_{i,k,t}^{(q)},$ (5)
where $\boldsymbol{w}_{k}$ is the coding vector at the $k$-th block. Collecting the received signals during $T$ time slots at a given frame $i$ and block $k$, we get
$\boldsymbol{Y}_{i,k}$ $=\sum\limits_{q=1}^{Q}\boldsymbol{H}^{(q)}\boldsymbol{S}_{k}^{(q)}\boldsymbol{G}^{(q)}_{i}\text{diag}(\boldsymbol{w}_{k})\boldsymbol{X}^{\text{T}}+\boldsymbol{Z}_{i,k}$ (6)
$=\Big(\sum\limits_{q=1}^{Q}\boldsymbol{H}^{(q)}\boldsymbol{S}_{k}^{(q)}\boldsymbol{G}^{(q)}_{i}\Big)\text{diag}(\boldsymbol{w}_{k})\boldsymbol{X}^{\text{T}}+\boldsymbol{Z}_{i,k}\in\mathbb{C}^{M_{r}\times T},$
where $\boldsymbol{X}=\left[\boldsymbol{x}_{1},\ldots,\boldsymbol{x}_{T}\right]^{\text{T}}\in\mathbb{C}^{T\times M_{T}}$ denotes the data symbol matrix, and $\boldsymbol{Z}_{i,k}\in\mathbb{C}^{M_{r}\times T}$ is the AWGN term [^3] Form this point of the document we suppressed the noise term for simplicity. However, in the simulations, the noise is included..
@@ II-A Scattering Matrix Design
The design of the BD-RIS scattering matrix used for training is fundamental to the uniqueness of the separate estimates of the involved communication channels. A baseline design for an LS channel estimation was first proposed in [B_Clerckc_CE_TSP_24, Clerckx_CAMSAP_2023], accounting for the physical constraints of the BD-RIS while ensuring energy conservation and the unitary property. Here, we consider an alternative scattering matrix design tailored to the proposed semi-blind channel and symbol estimation approach. Starting from the fundamental signal model in (6), we propose designing the BD-RIS matrix as the product of a fixed (static) scattering matrix $\boldsymbol{S}_{0}$ and a time-varying block-dependent diagonal rotation [^4] In contrast to the traditional BD-RIS modeling in [B_Clerckc_CE_TSP_24], where the scattering matrix is constant, the work in [Sokal_BD_RIS_2024] proposes varying the scattering matrix over time. However, this approach implies dynamically changing physical connections, which can be complex to implement. In our method, this factorization simplifies analysis and optimization by decoupling the immutable physical structure from the reconfigurable electronic controls. Specifically, $\boldsymbol{S}_{0}$ can represent the surface’s static, inherent physical properties—such as inter-element coupling and mutual impedance, which define the complex wave-propagation pathways. In contrast, $\boldsymbol{P}$ can represent the active control that applies individual phase shifts to dynamically steer the overall response (similar to the traditional RIS)., such that $\boldsymbol{S}_{k}^{(q)}=\boldsymbol{S}_{0}\text{diag}(\boldsymbol{\overline{p}}_{k}^{(q)})$ where $\boldsymbol{\overline{p}}_{k}^{(q)}$ is the vector containing the phase rotation angles associated with the $q$-th group and the $k$-th block. Different from the state-of-the-art [B_Clerckc_CE_TSP_24] and [Sokal_BD_RIS_2024], the proposed design has a lower complex implementation since only the diagonal term $\boldsymbol{\overline{p}}_{k}^{(q)}$ changes per group. The proposed design also maintains the unitary property of the scattering matrix, as in [B_Clerckc_CE_TSP_24], i.e.,
$(\boldsymbol{S}_{k}^{(q)})^{\text{H}}\boldsymbol{S}_{k}^{(q)}=\boldsymbol{S}_{k}^{(q)}(\boldsymbol{S}_{k}^{(q)})^{\text{H}}=\boldsymbol{I}_{\overline{N}}\;\;\;\forall q.$ (7)
@@
To meet this constraint, the static component $\boldsymbol{S}_{0}$ can be chosen as any unitary/orthogonal matrix (e.g., a truncated discrete Fourier transform (DFT) or Hadamard matrix). On the other hand, to ensure the identity in (7), the following condition must be satisfied:
$\text{diag}(\boldsymbol{\overline{p}}_{k}^{(q)})^{\text{H}}\text{diag}(\boldsymbol{\overline{p}}_{k}^{(q)})=\text{diag}(\boldsymbol{\overline{p}}_{k}^{(q)})\text{diag}(\boldsymbol{\overline{p}}_{k}^{(q)})^{\text{H}}=\boldsymbol{I}_{\overline{N}}\;\;\forall q.$ (8)
@@
Define $\overline{\boldsymbol{P}}^{(q)}=[\boldsymbol{\overline{p}}_{1}^{(q)},\ldots,\boldsymbol{\overline{p}}_{K}^{(q)}]^{\text{T}}\in\mathbb{C}^{K\times\overline{N}}$, the diagonal rotation component can be chosen to meet the following condition
$\Big(\overline{\boldsymbol{P}}^{(\text{q})}\Big)^{\ast}\odot\overline{\boldsymbol{P}}^{(q)}=\overline{\boldsymbol{P}}^{(q)}\odot\Big(\overline{\boldsymbol{P}}^{(\text{q})}\Big)^{\ast}=\boldsymbol{1}_{K\times\overline{N}}\;\;\;\forall q.$ (9)
@@
$\left|\overline{\boldsymbol{P}}^{(\text{q})}_{k\overline{n}}\right|=1,\;\forall\;k=1,\,\dots,\,K\;\text{and}\;\forall\;\overline{n}=1,\,\dots,\,\overline{N},$ (10)
which means that the entries of the rotation matrix must belong to the complex unitary circle, for example, $\overline{\boldsymbol{P}}^{(\text{q})}_{k\overline{n}}=e^{i\theta_{k\overline{n}}}$ with $\theta_{k\overline{n}}\in[0,\,2\pi)$.
@@ II-B Received Signal Model
$\boldsymbol{Y}_{i,k}$ $=\Big(\sum_{q=1}^{Q}\boldsymbol{H}^{(q)}\boldsymbol{S}_{0}^{(q)}\text{diag}(\boldsymbol{\overline{p}}_{k}^{(q)})\boldsymbol{G}^{(q)}_{i}\Big)\text{diag}(\boldsymbol{w}_{k})\boldsymbol{X}^{\text{T}},$ (11)
@@
$\boldsymbol{Y}_{i,k}$ $=\Big[\boldsymbol{H}^{(1)}\boldsymbol{S}_{0}^{(1)}\dots\boldsymbol{H}^{(Q)}\boldsymbol{S}_{0}^{(Q)}\Big]\text{diag}(\boldsymbol{p}_{k})\boldsymbol{G}_{i}\text{diag}(\boldsymbol{w}_{k})\boldsymbol{X}^{\text{T}}$ (12)
$=\Big[\boldsymbol{H}^{(1)}\dots\boldsymbol{H}^{(Q)}\Big]\boldsymbol{S}\text{diag}(\boldsymbol{p}_{k})\boldsymbol{G}_{i}\text{diag}(\boldsymbol{w}_{k})\boldsymbol{X}^{\text{T}}$
$=\boldsymbol{H}\boldsymbol{S}\text{diag}(\boldsymbol{p}_{k})\boldsymbol{G}_{i}\text{diag}(\boldsymbol{w}_{k})\boldsymbol{X}^{\text{T}}$
@@
$\boldsymbol{Y}_{i,k}=\boldsymbol{H}\boldsymbol{S}\boldsymbol{D}_{k}(\boldsymbol{P})\boldsymbol{G}_{i}\boldsymbol{D}_{k}(\boldsymbol{W})\boldsymbol{X}^{\text{T}},$ (13)
where $\boldsymbol{G}_{i}=\left[(\boldsymbol{G}_{i}^{(1)})^{\text{T}}\;\dots\;(\boldsymbol{G}_{i}^{(Q)})^{\text{T}}\right]^{\text{T}}\in\mathbb{C}^{N\times M_{T}}$ and $\boldsymbol{H}=\left[\boldsymbol{H}^{(1)}\;\dots\;\boldsymbol{H}^{(Q)}\right]\in\mathbb{C}^{M_{R}\times N}$ are the two involved channel matrices collecting the contributions from all groups. Let us define $\boldsymbol{p}_{k}\doteq[\boldsymbol{p}^{(1)\text{T}}_{k},\ldots,\boldsymbol{p}^{(Q)\text{T}}_{k}]^{\text{T}}\in\mathbb{C}^{N\times 1}$ and $\boldsymbol{S}\doteq\text{bdiag}(\boldsymbol{S}_{0}^{(1)},\ldots\boldsymbol{S}_{0}^{(Q)})\in\mathbb{C}^{N\times N}$, and note that $\boldsymbol{S}\boldsymbol{S}^{\text{H}}=\boldsymbol{I}_{N}$ since $\boldsymbol{S}_{0}^{(q)}\boldsymbol{S}_{0}^{(q)\text{H}}=\boldsymbol{I}_{\overline{N}}$. Additionally, define $\boldsymbol{P}=[\boldsymbol{p}_{1},\ldots,\boldsymbol{p}_{K}]^{\text{T}}\in\mathbb{C}^{K\times N}$, with $\boldsymbol{P}^{\text{H}}\odot\boldsymbol{P}=\boldsymbol{P}\odot\boldsymbol{P}^{\text{H}}=\boldsymbol{1}_{N}$ and $\boldsymbol{W}\doteq[\boldsymbol{w}_{1},\ldots,\boldsymbol{w}_{K}]^{\text{T}}\in\mathbb{C}^{K\times M_{T}}$ as the BD-RIS phase rotation matrix and transmit coding matrix, respectively.
The noiseless signal model in (13) corresponds to a PARATUCK-(2,4) decomposition of the fourth-order tensor $\boldsymbol{\mathcal{Y}}\in\mathbb{C}^{M_{R}\times T\times K\times I}$, which belongs to the family of PARATUCK-type decompositions [Andre_EURASIP_2014]. As a hybrid of the PARAFAC and TUCKER decompositions, it combines the uniqueness of the former with the flexibility of the latter.
@@ III-A PARAFAC-Based Receiver
The received signal model in (13) involves three unknown system matrices, namely, the static channel matrix $\boldsymbol{H}$ linking the BD-RIS and the BS, the block-dependent channel matrix $\boldsymbol{G}_{i}$ linking the UT and BD-RIS, and the symbol matrix $\boldsymbol{X}$.
@@
$\boldsymbol{y}_{i,k}=\text{vec}(\boldsymbol{Y}_{i,k})=\Big(\boldsymbol{X}\otimes\boldsymbol{HS}\Big)\Big(\boldsymbol{D}_{k}(\boldsymbol{W})\otimes\boldsymbol{D}_{k}(\boldsymbol{P})\Big)\boldsymbol{g}_{i},$ (14)
@@
$\text{vec}(\boldsymbol{Y}_{i,k})=\Big(\boldsymbol{X}\otimes\boldsymbol{HS}\Big)\text{diag}(\boldsymbol{g}_{i})\Big(\boldsymbol{W}_{k\cdot}^{\text{T}}\otimes\boldsymbol{P}_{k\cdot}^{\text{T}}\Big)\;.$ (15)
Then, collecting the signal from all of the $K$ data blocks $\boldsymbol{Y}_{i}=\Big[\text{vec}(\boldsymbol{Y}_{i,1})\;\text{vec}(\boldsymbol{Y}_{i,2})\;\dots\;\text{vec}(\boldsymbol{Y}_{i,K})\Big]$, we have the received signal at frame $i$ as
$\boldsymbol{Y}_{i}$ $=\Big[\boldsymbol{y}_{i,1}\;\boldsymbol{y}_{i,2}\;\dots\;\boldsymbol{y}_{i,K}\Big]$ (16)
$=\Big(\boldsymbol{X}\otimes\boldsymbol{HS}\Big)\text{diag}(\boldsymbol{g}_{i})\Big[\boldsymbol{W}_{1\cdot}^{\text{T}}\otimes\boldsymbol{P}_{1\cdot}^{\text{T}}\;\dots\;\boldsymbol{W}_{K\cdot}^{\text{T}}\otimes\boldsymbol{P}_{K\cdot}^{\text{T}}\Big]\;$
$=\Big(\boldsymbol{X}\otimes\boldsymbol{HS}\Big)\text{diag}(\boldsymbol{g}_{i})\Big(\boldsymbol{W}^{\text{T}}\diamond\boldsymbol{P}^{\text{T}}\Big)\;\in\mathbb{C}^{M_{R}T\times K}.$
Indeed, (16), is a $3$-mode unfoldding transposed $\boldsymbol{\mathcal{Y}}_{i}]_{(3)}^{\text{T}}$ of the tensor $\boldsymbol{\mathcal{Y}}_{i}\in\mathbb{C}^{M_{R}\times T\times K}$ that fits a third-order TUCKER tensor model whose the $n$-mode product is given as
$\boldsymbol{\mathcal{Y}}_{i}=\boldsymbol{\mathcal{T}}_{\boldsymbol{G}_{i}}\times_{1}\boldsymbol{HS}\times_{2}\boldsymbol{X}\times_{3}\Big(\boldsymbol{W}^{\text{T}}\diamond\boldsymbol{P}^{\text{T}}\Big)^{\text{T}},$ (17)
where the $3$-mode unfolding of the tensor $\boldsymbol{\mathcal{T}}_{\boldsymbol{G}_{i}}]_{(3)}$ coincides with $\text{diag}(\boldsymbol{g}_{i})$, that is, $\boldsymbol{\mathcal{T}}_{\boldsymbol{G}_{i}}]_{(3)}=\text{diag}(\boldsymbol{g}_{i})$. Defining
$\boldsymbol{\Omega}\doteq\Big(\boldsymbol{X}\otimes\boldsymbol{HS}\Big)\in\mathbb{C}^{TM_{R}\times M_{T}N},$ (18)
@@
$\boldsymbol{\Psi}^{\text{T}}\doteq\Big(\boldsymbol{W}^{\text{T}}\diamond\boldsymbol{P}^{\text{T}}\Big)\in\mathbb{C}^{M_{T}N\times K},$ (19)
@@
$\boldsymbol{\mathcal{Y}}_{i}]_{(3)}^{\text{T}}=\boldsymbol{\Omega}[\boldsymbol{\mathcal{T}}_{\boldsymbol{G}_{i}}]_{(3)}\boldsymbol{\Psi}^{\text{T}}\in\mathbb{C}^{TM_{R}\times K}\;.$ (20)
Let $\boldsymbol{\mathcal{Z}}\in\mathbb{C}^{TM_{R}\times K\times I}$ be a third order tensor which is defined as
$\boldsymbol{\mathcal{Z}}\doteq[\boldsymbol{\mathcal{Y}}_{1}]_{(3)}^{\text{T}}\sqcup_{3}[\boldsymbol{\mathcal{Y}}_{2}]_{(3)}^{\text{T}}\dots\sqcup_{3}[\boldsymbol{\mathcal{Y}}_{I}]_{(3)}^{\text{T}},$ (21)

### v2

@@
Gilderlan Tavares de Araújo, André L. F. de Almeida, Bruno Sokal, Gabor Fodor, Paulo R. B Gomes The authors thank the partial support of FUNCAP under grant ITR-0214-00041.01.00/23, and the National Institute of Science and Technology (INCT-Signals) sponsored by Brazil’s National Council for Scientific and Technological Development (CNPq) under grant 406517/2022-3.Gilderlan T. de Araújo and Paulo. R. B. Gomes are with the Federal Institute of Ceará, e-mails: {gilerlan.tavares,gomes.paulo}@ifce.edu.br.André L. F. de Almeida and Bruno Sokal are with the Teleinformatics Department, Federal University of Ceará, Fortaleza-CE, e-mail: {andre,brunosokal}@gtel.ufc.br.G. Fodor was funded by the Swedish Strategic Foundation Future Software Systems Project SAICOM, grant no: FUS21-0004
@@ I Introduction
$\textrm{vec}(\bm{ABC})=(\bm{C}^{\textrm{T}}\otimes\bm{A})\textrm{vec}(\bm{B}).$ (1)
$\textrm{diag}(\bm{a})\bm{b}=\textrm{diag}(\bm{b})\bm{a}.$ (2)
If $\bm{B}$ is a diagonal matrix, we have:
$\textrm{vec}(\bm{ABC})=(\bm{C}^{\textrm{T}}\diamond\bm{A})\textrm{vecd}(\bm{B}).$ (3)
@@ II System Model
$\bm{y}_{t}=\sum_{q=1}^{Q}\bm{H}^{(q)}\bm{S}^{(q)}\bm{G}^{(q)}\bm{x}_{t}+\bm{z}_{t}^{(q)},$ (4)
where $\bm{H}^{(q)}\in\mathbb{C}^{M_{R}\times\overline{N}}$ and $\bm{G}^{(q)}\in\mathbb{C}^{\overline{N}\times M_{T}}$ represent the RIS-BS and UT-RIS channels, respectively, while $\bm{S}^{(q)}\in\mathbb{C}^{\overline{N}\times\overline{N}}$ denotes the (non-diagonal) scattering matrix associated with the $q$-th group. Additionally, $\bm{x}_{t}\in\mathbb{C}^{M_{T}\times 1}$ represents the data symbol vector, while $\bm{z}_{t}^{(q)}$ corresponds to the additive white Gaussian noise (AWGN) term.
@@
$\bm{y}_{i,k,t}=\sum_{q=1}^{Q}\bm{H}^{(q)}\bm{S}_{k}^{(q)}\bm{G}^{(q)}_{i}\text{diag}(\bm{w}_{k})\bm{x}_{t}+\bm{z}_{i,k,t}^{(q)},$ (5)
where $\bm{w}_{k}$ is the coding vector at the $k$-th block. Collecting the received signals during $T$ time slots at a given frame $i$ and block $k$, we get
$\bm{Y}_{i,k}$ $=\sum\limits_{q=1}^{Q}\bm{H}^{(q)}\bm{S}_{k}^{(q)}\bm{G}^{(q)}_{i}\text{diag}(\bm{w}_{k})\bm{X}^{\text{T}}+\bm{Z}_{i,k}$ (6)
$=\Big(\sum\limits_{q=1}^{Q}\bm{H}^{(q)}\bm{S}_{k}^{(q)}\bm{G}^{(q)}_{i}\Big)\text{diag}(\bm{w}_{k})\bm{X}^{\text{T}}+\bm{Z}_{i,k}\in\mathbb{C}^{M_{r}\times T},$
where $\bm{X}=\left[\bm{x}_{1},\ldots,\bm{x}_{T}\right]^{\text{T}}\in\mathbb{C}^{T\times M_{T}}$ denotes the data symbol matrix, and $\bm{Z}_{i,k}\in\mathbb{C}^{M_{r}\times T}$ is the AWGN term [^3] Form this point of the document we suppressed the noise term for simplicity. However, in the simulations, the noise is included..
@@ II-A Scattering Matrix Design
The design of the BD-RIS scattering matrix used for training is fundamental to the uniqueness of the separate estimates of the involved communication channels. A baseline design for an LS channel estimation was first proposed in [B_Clerckc_CE_TSP_24, Clerckx_CAMSAP_2023], accounting for the physical constraints of the BD-RIS while ensuring energy conservation and the unitary property. Here, we consider an alternative scattering matrix design tailored to the proposed semi-blind channel and symbol estimation approach. Starting from the fundamental signal model in (6), we propose designing the BD-RIS matrix as the product of a fixed (static) scattering matrix $\bm{S}_{0}$ and a time-varying block-dependent diagonal rotation [^4] In contrast to the traditional BD-RIS modeling in [B_Clerckc_CE_TSP_24], where the scattering matrix is constant, the work in [Sokal_BD_RIS_2024] proposes varying the scattering matrix over time. However, this approach implies dynamically changing physical connections, which can be complex to implement. In our method, this factorization simplifies analysis and optimization by decoupling the immutable physical structure from the reconfigurable electronic controls. Specifically, $\bm{S}_{0}$ can represent the surface’s static, inherent physical properties—such as inter-element coupling and mutual impedance, which define the complex wave-propagation pathways. In contrast, $\bm{P}$ can represent the active control that applies individual phase shifts to dynamically steer the overall response (similar to the traditional RIS)., such that $\bm{S}_{k}^{(q)}=\bm{S}_{0}\text{diag}(\bm{\overline{p}}_{k}^{(q)})$ where $\bm{\overline{p}}_{k}^{(q)}$ is the vector containing the phase rotation angles associated with the $q$-th group and the $k$-th block. Different from the state-of-the-art [B_Clerckc_CE_TSP_24] and [Sokal_BD_RIS_2024], the proposed design has a lower complex implementation since only the diagonal term $\bm{\overline{p}}_{k}^{(q)}$ changes per group. The proposed design also maintains the unitary property of the scattering matrix, as in [B_Clerckc_CE_TSP_24], i.e.,
$(\bm{S}_{k}^{(q)})^{\text{H}}\bm{S}_{k}^{(q)}=\bm{S}_{k}^{(q)}(\bm{S}_{k}^{(q)})^{\text{H}}=\bm{I}_{\overline{N}}\;\;\;\forall q.$ (7)
@@
To meet this constraint, the static component $\bm{S}_{0}$ can be chosen as any unitary/orthogonal matrix (e.g., a truncated discrete Fourier transform (DFT) or Hadamard matrix). On the other hand, to ensure the identity in (7), the following condition must be satisfied:
$\text{diag}(\bm{\overline{p}}_{k}^{(q)})^{\text{H}}\text{diag}(\bm{\overline{p}}_{k}^{(q)})=\text{diag}(\bm{\overline{p}}_{k}^{(q)})\text{diag}(\bm{\overline{p}}_{k}^{(q)})^{\text{H}}=\bm{I}_{\overline{N}}\;\;\forall q.$ (8)
@@
Define $\overline{\bm{P}}^{(q)}=[\bm{\overline{p}}_{1}^{(q)},\ldots,\bm{\overline{p}}_{K}^{(q)}]^{\text{T}}\in\mathbb{C}^{K\times\overline{N}}$, the diagonal rotation component can be chosen to meet the following condition
$\Big(\overline{\bm{P}}^{(\text{q})}\Big)^{\ast}\odot\overline{\bm{P}}^{(q)}=\overline{\bm{P}}^{(q)}\odot\Big(\overline{\bm{P}}^{(\text{q})}\Big)^{\ast}=\bm{1}_{K\times\overline{N}}\;\;\;\forall q.$ (9)
@@
$\left|\overline{\bm{P}}^{(\text{q})}_{k\overline{n}}\right|=1,\;\forall\;k=1,\,\dots,\,K\;\text{and}\;\forall\;\overline{n}=1,\,\dots,\,\overline{N},$ (10)
which means that the entries of the rotation matrix must belong to the complex unitary circle, for example, $\overline{\bm{P}}^{(\text{q})}_{k\overline{n}}=e^{i\theta_{k\overline{n}}}$ with $\theta_{k\overline{n}}\in[0,\,2\pi)$.
@@ II-B Received Signal Model
$\bm{Y}_{i,k}$ $=\Big(\sum_{q=1}^{Q}\bm{H}^{(q)}\bm{S}_{0}^{(q)}\text{diag}(\bm{\overline{p}}_{k}^{(q)})\bm{G}^{(q)}_{i}\Big)\text{diag}(\bm{w}_{k})\bm{X}^{\text{T}},$ (11)
@@
$\bm{Y}_{i,k}$ $=\Big[\bm{H}^{(1)}\bm{S}_{0}^{(1)}\dots\bm{H}^{(Q)}\bm{S}_{0}^{(Q)}\Big]\text{diag}(\bm{p}_{k})\bm{G}_{i}\text{diag}(\bm{w}_{k})\bm{X}^{\text{T}}$ (12)
$=\Big[\bm{H}^{(1)}\dots\bm{H}^{(Q)}\Big]\bm{S}\text{diag}(\bm{p}_{k})\bm{G}_{i}\text{diag}(\bm{w}_{k})\bm{X}^{\text{T}}$
$=\bm{H}\bm{S}\text{diag}(\bm{p}_{k})\bm{G}_{i}\text{diag}(\bm{w}_{k})\bm{X}^{\text{T}}$
@@
$\bm{Y}_{i,k}=\bm{H}\bm{S}\bm{D}_{k}(\bm{P})\bm{G}_{i}\bm{D}_{k}(\bm{W})\bm{X}^{\text{T}},$ (13)
where $\bm{G}_{i}=\left[(\bm{G}_{i}^{(1)})^{\text{T}}\;\dots\;(\bm{G}_{i}^{(Q)})^{\text{T}}\right]^{\text{T}}\in\mathbb{C}^{N\times M_{T}}$ and $\bm{H}=\left[\bm{H}^{(1)}\;\dots\;\bm{H}^{(Q)}\right]\in\mathbb{C}^{M_{R}\times N}$ are the two involved channel matrices collecting the contributions from all groups. Let us define $\bm{p}_{k}\doteq[\bm{p}^{(1)\text{T}}_{k},\ldots,\bm{p}^{(Q)\text{T}}_{k}]^{\text{T}}\in\mathbb{C}^{N\times 1}$ and $\bm{S}\doteq\text{bdiag}(\bm{S}_{0}^{(1)},\ldots\bm{S}_{0}^{(Q)})\in\mathbb{C}^{N\times N}$, and note that $\bm{S}\bm{S}^{\text{H}}=\bm{I}_{N}$ since $\bm{S}_{0}^{(q)}\bm{S}_{0}^{(q)\text{H}}=\bm{I}_{\overline{N}}$. Additionally, define $\bm{P}=[\bm{p}_{1},\ldots,\bm{p}_{K}]^{\text{T}}\in\mathbb{C}^{K\times N}$, with $\bm{P}^{\text{H}}\odot\bm{P}=\bm{P}\odot\bm{P}^{\text{H}}=\bm{1}_{N}$ and $\bm{W}\doteq[\bm{w}_{1},\ldots,\bm{w}_{K}]^{\text{T}}\in\mathbb{C}^{K\times M_{T}}$ as the BD-RIS phase rotation matrix and transmit coding matrix, respectively.
The noiseless signal model in (13) corresponds to a PARATUCK-(2,4) decomposition of the fourth-order tensor $\bm{\mathcal{Y}}\in\mathbb{C}^{M_{R}\times T\times K\times I}$, which belongs to the family of PARATUCK-type decompositions [Andre_EURASIP_2014]. As a hybrid of the PARAFAC and TUCKER decompositions, it combines the uniqueness of the former with the flexibility of the latter.
@@ III-A PARAFAC-Based Receiver
The received signal model in (13) involves three unknown system matrices, namely, the static channel matrix $\bm{H}$ linking the BD-RIS and the BS, the block-dependent channel matrix $\bm{G}_{i}$ linking the UT and BD-RIS, and the symbol matrix $\bm{X}$.
@@
$\bm{y}_{i,k}=\text{vec}(\bm{Y}_{i,k})=\Big(\bm{X}\otimes\bm{HS}\Big)\Big(\bm{D}_{k}(\bm{W})\otimes\bm{D}_{k}(\bm{P})\Big)\bm{g}_{i},$ (14)
@@
$\text{vec}(\bm{Y}_{i,k})=\Big(\bm{X}\otimes\bm{HS}\Big)\text{diag}(\bm{g}_{i})\Big(\bm{W}_{k\cdot}^{\text{T}}\otimes\bm{P}_{k\cdot}^{\text{T}}\Big)\;.$ (15)
Then, collecting the signal from all of the $K$ data blocks $\bm{Y}_{i}=\Big[\text{vec}(\bm{Y}_{i,1})\;\text{vec}(\bm{Y}_{i,2})\;\dots\;\text{vec}(\bm{Y}_{i,K})\Big]$, we have the received signal at frame $i$ as
$\bm{Y}_{i}$ $=\Big[\bm{y}_{i,1}\;\bm{y}_{i,2}\;\dots\;\bm{y}_{i,K}\Big]$ (16)
$=\Big(\bm{X}\otimes\bm{HS}\Big)\text{diag}(\bm{g}_{i})\Big[\bm{W}_{1\cdot}^{\text{T}}\otimes\bm{P}_{1\cdot}^{\text{T}}\;\dots\;\bm{W}_{K\cdot}^{\text{T}}\otimes\bm{P}_{K\cdot}^{\text{T}}\Big]\;$
$=\Big(\bm{X}\otimes\bm{HS}\Big)\text{diag}(\bm{g}_{i})\Big(\bm{W}^{\text{T}}\diamond\bm{P}^{\text{T}}\Big)\;\in\mathbb{C}^{M_{R}T\times K}.$
Indeed, (16), is a $3$-mode unfoldding transposed $\bm{\mathcal{Y}}_{i}]_{(3)}^{\text{T}}$ of the tensor $\bm{\mathcal{Y}}_{i}\in\mathbb{C}^{M_{R}\times T\times K}$ that fits a third-order TUCKER tensor model whose the $n$-mode product is given as
$\bm{\mathcal{Y}}_{i}=\bm{\mathcal{T}}_{\bm{G}_{i}}\times_{1}\bm{HS}\times_{2}\bm{X}\times_{3}\Big(\bm{W}^{\text{T}}\diamond\bm{P}^{\text{T}}\Big)^{\text{T}},$ (17)
where the $3$-mode unfolding of the tensor $\bm{\mathcal{T}}_{\bm{G}_{i}}]_{(3)}$ coincides with $\text{diag}(\bm{g}_{i})$, that is, $\bm{\mathcal{T}}_{\bm{G}_{i}}]_{(3)}=\text{diag}(\bm{g}_{i})$. Defining
$\bm{\Omega}\doteq\Big(\bm{X}\otimes\bm{HS}\Big)\in\mathbb{C}^{TM_{R}\times M_{T}N},$ (18)
@@
$\bm{\Psi}^{\text{T}}\doteq\Big(\bm{W}^{\text{T}}\diamond\bm{P}^{\text{T}}\Big)\in\mathbb{C}^{M_{T}N\times K},$ (19)
@@
$\bm{\mathcal{Y}}_{i}]_{(3)}^{\text{T}}=\bm{\Omega}[\bm{\mathcal{T}}_{\bm{G}_{i}}]_{(3)}\bm{\Psi}^{\text{T}}\in\mathbb{C}^{TM_{R}\times K}\;.$ (20)
Let $\bm{\mathcal{Z}}\in\mathbb{C}^{TM_{R}\times K\times I}$ be a third order tensor which is defined as
$\bm{\mathcal{Z}}\doteq[\bm{\mathcal{Y}}_{1}]_{(3)}^{\text{T}}\sqcup_{3}[\bm{\mathcal{Y}}_{2}]_{(3)}^{\text{T}}\dots\sqcup_{3}[\bm{\mathcal{Y}}_{I}]_{(3)}^{\text{T}},$ (21)